OpenStudy (anonymous):

5(x - 5) = 4(x - 1) + 3x.

OpenStudy (anonymous):

\[5x -25 = 4x-4 +3x\]

OpenStudy (anonymous):

5x-4x-3x = 25-4

OpenStudy (anonymous):

-2x = 21 x=21/-2

OpenStudy (anonymous):

i got no -2 on my answer list

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

did you put the question in right?